A) An economy produces three goods: organ, apple and organ juice. It is assumed that half of the organ are bought and consumed as final good; the other half is used to produce organ juice. 


In 2019, 20 KGs of organ are produced at Rs.50/kg, 40 KGs of apple are produced at 100/kg, and 25 Glasses of organ juice are produced at Rs.200/glass.


In 2020, 40 KGs of organ are produced at Rs.80/kg, 60 KGs of apple are produced at 110/kg, and 30 glasses of organ juice are produced at Rs.195/glass.


  Find the following:


(i) Assuming 2019 is the base year; calculate the real and nominal GDP in 2019 and 2020.


B) Analyze how much Pakistan’s GDP and each of its components is affected by the following transactions? Explain your answers


(i) Atiya spends Rs.1000 to purchase dinner at the finest restaurant in Karachi


(ii)Your mother spend weekend cleaning your house


(iii)Dalda Banspati Limited Pakistan builds Rs.50 million worth of cooking oil, but consumers only buy Rs. 45 million worth of them.



1
Expert's answer
2021-01-26T08:49:47-0500

(i) The real and nominal GDP in 2019 are:

Real GDP = Nominal GDP = 20×50 + 40×100 + 25×200 = 10,000.

In 2020 nominal GDP = 40×80 + 60×110 + 30×195 = 15,650.

Real GDP = 40×50 + 60×100 + 30×200 = 14,000.

B)

(i) If Atiya spends Rs.1000 to purchase dinner at the finest restaurant in Karachi, then consumption will increase by Rs. 1000.

(ii) If your mother spends weekend cleaning your house, then GDP will not be affected.

(iii) If Dalda Banspati Limited Pakistan builds Rs.50 million worth of cooking oil, but consumers only buy Rs. 45 million worth of them, then consumption will decrease by Rs. 5 million.